OpenStudy (anonymous):

Is g(x) differentiable at 0? Is g(x) differentiable at 2?...

OpenStudy (anonymous):

2x if x is less than or equal to 0 2-x^(2) if 0<x<2 2-x if x is greater than or equal to 2

OpenStudy (anonymous):

Piecewise function :) Thanks!

jimthompson5910 (jim_thompson5910):

Hint: if g ' (x) is continuous at x = c, then g(x) is differentiable at x = c

OpenStudy (anonymous):

Is it continuous at x=2?

jimthompson5910 (jim_thompson5910):

if y = 2x, then y ' = ???

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

if y = 2 - x^2, then y ' = ???

OpenStudy (anonymous):

2-2x

jimthompson5910 (jim_thompson5910):

oh also, I should have stated that g(x) needs to be continuous at x = c to be differentiable at x = c (along with the hint I provided)

jimthompson5910 (jim_thompson5910):

no just -2x

jimthompson5910 (jim_thompson5910):

if y = 2 - x, then y ' = ???

OpenStudy (anonymous):

but wouldn't it be 2x-x^(2)= 2(1)-2x^(2-1)

OpenStudy (anonymous):

and the last one is -1?

jimthompson5910 (jim_thompson5910):

wait, the original is 2x - x^2 ? or 2 - x^2 ?

OpenStudy (anonymous):

Oh, I forgot the x.

OpenStudy (anonymous):

Sorry about that.

jimthompson5910 (jim_thompson5910):

ok so the second piece is 2x - x^2 correct?

jimthompson5910 (jim_thompson5910):

just want to confirm

OpenStudy (anonymous):

Yes.

jimthompson5910 (jim_thompson5910):

ok one moment

jimthompson5910 (jim_thompson5910):

I'm going to let p(x) be the piecewise function f(x) will be the first piece (I've graphed in red) g(x) is the second piece (graphed in blue) h(x) is the third piece (in green) see attached Notice how the different pieces connect. So this function is definitely continuous at x = 0 and x = 2. It is continuous everywhere else as well.

jimthompson5910 (jim_thompson5910):

now if we derive p(x), we have to derive everything that makes up p(x) so we derive f(x) to get f ' (x), g to get g ', etc etc this is shown in the next attachment notice how p ' (x) is continuous at x = 0 (the red and blue pieces connect); however, p ' (x) is NOT continuous at x = 2. There's a jump from (2,-2) on the blue piece to (2,-1) on the green piece. Because of this discontinuity on p ' (x) at x = 2, this means p(x) is NOT differentiable at x = 2

jimthompson5910 (jim_thompson5910):

If you look back at the first attachment http://assets.openstudy.com/updates/attachments/54484a4be4b0f84976800a4f-jim_thompson5910-1414025861146-1.png you'll notice that at x = 2, there's a sharp point. This is usually a good indication the function isn't differentiable here. However, it's always a good idea to derive and find out.

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome